# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

### v1

@@ Appendix A Proving Inequalities Between PDP and ICE-based global sensitivity metrics
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j}),$
@@
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})+h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]=I_{pdp}(x_{j})$.
@@
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$. Moreover, equality holds if $h$ is a nonnegative function.
@@
If $f(x_{j},\boldsymbol{x}_{C})=g_{1}(x_{j})h_{1}(\boldsymbol{x}_{C})+\ldots+g_{k}(x_{j})h_{k}(\boldsymbol{x}_{C})$, where $g_{1},\ldots,g_{k}$ are functions of $x_{j}$ only and $h_{1},\ldots,h_{k}$ are functions of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$.
@@
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{H(\boldsymbol{x}_{C})^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))H(\boldsymbol{x}_{C})}\right].$
@@
If $f$ is a polynomial in $x_{1},\ldots,x_{m}$, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$ for any $j=1,\ldots,m$.
@@
We conjecture that the inequality also holds for more general functions, though an investigation of this claim will be left for future work.

### v3

@@ Appendix A Proving Inequalities Between PDP and ICE-based global sensitivity metrics
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j}),$
@@
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})+h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=I_{pdp}(x_{j})$.
@@
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$. Moreover, equality holds if $h$ is a nonnegative function.
@@
If $f(x_{j},\boldsymbol{x}_{C})=g_{1}(x_{j})h_{1}(\boldsymbol{x}_{C})+\ldots+g_{k}(x_{j})h_{k}(\boldsymbol{x}_{C})$, where $g_{1},\ldots,g_{k}$ are functions of $x_{j}$ only and $h_{1},\ldots,h_{k}$ are functions of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$.
@@
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{H(\boldsymbol{x}_{C})^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))H(\boldsymbol{x}_{C})}\right].$
@@
If $f$ is a polynomial in $x_{1},\ldots,x_{m}$, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$ for any $j=1,\ldots,m$.
@@
Various surrogate and machine-learning models admit such a structure, either exactly or approximately. Examples include support vector regression models with polynomial kernels, as well as expressions obtained via genetic programming when the search space is constrained to enforce additive or multiplicative separability. However, for more complex models, such as feedforward neural networks with multiple nonlinear hidden layers, the above proof is not directly applicable due to the absence of explicit additive or multiplicative separability. We conjecture that the inequality also holds for more general functions, though an investigation of this claim will be left for future work.